OpenStudy (anonymous):

What is an equation of the line that passes through the points (-4, 3) and (-1, 6)?

OpenStudy (phi):

can you find the slope between those two points ?

OpenStudy (anonymous):

Point A (-4, 3) Point B (-1, 6) First find the slope of Line AB: s = (ya - yB)/(xA - xB) = (6 - 3)/(-1 + 4) = 1. The equation of the line y - x + b. Find b by writing that this line passes at A (-4, 3): (3) = (-4) + b --> b = ... The answer is y = x + ....